OpenStudy (anonymous):

how to find the percent of change.Round to the nearest tenth, If necessary. Then state whether the percent of change is an increase or decrease? from 12 m to 18 m

OpenStudy (anonymous):

To find the change is to subtract 12m from 18m

OpenStudy (anonymous):

To find the percentage change: The answer you get divide it by the distance that you started out with then multiply that by 100 to put your answer in percentage.

OpenStudy (anonymous):

Can you try this and tell me what you get?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Are you having problems?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Subtract 12 from 18: 18-12 = 6m Divide 6 by the distance you started with (12): 6/12 =0.5(100) to put ans in %

OpenStudy (anonymous):

you should get 50%

OpenStudy (anonymous):

Now to determine if there was a decrease or increase. What distance did we start out with?

OpenStudy (anonymous):

Are you still there?

OpenStudy (anonymous):

ya sorry

OpenStudy (anonymous):

an decrease?

OpenStudy (anonymous):

What distance did we start out from?

OpenStudy (anonymous):

Answer:You started at 12m and ended up at 18m that means that the change is increasing. Number on the right hand side is bigger, thus increase.

OpenStudy (anonymous):

If you were given another question where the distance moved from 18m to 12m then that means decreasing. (number on the right is the smallest)

OpenStudy (anonymous):

Hope this helps you with understanding this topic.